OpenStudy (anonymous):

i need to use the quadratic formula to solve xsquared-4x-7=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

easier to complete the square because middle term is even \[x^2-4x=7\] \[(x-2)^2=7+4=11\] \[x-2=\pm\sqrt{11}\] \[x=2\pm\sqrt{11}\]

OpenStudy (anonymous):

thanks